OpenStudy (zzelinski):

I know this is really simple, but I seem to have forgotten how to do this problem. -1/14-3/5 help please?

jimthompson5910 (jim_thompson5910):

the LCD is 14*5 = 70

jimthompson5910 (jim_thompson5910):

-1/14-3/5 you need to get every denominator equal to the LCD before you can subtract the fractions (-1/14)*(5/5)-3/5 -5/70-3/5 -5/70-(3/5)(14/14) -5/70-42/70 (-5 - 42)/70 -47/70 So -1/14-3/5 = -47/70

OpenStudy (zzelinski):

Thanks! I have a couple more questions if that's alright?

jimthompson5910 (jim_thompson5910):

ok just a few more though

OpenStudy (zzelinski):

7/8y-5/8=15/4

jimthompson5910 (jim_thompson5910):

y is in the denominator?

OpenStudy (zzelinski):

yeah

OpenStudy (zzelinski):

oh wait no its \[\frac{ 7 }{ 8} y \]

jimthompson5910 (jim_thompson5910):

ok start off by multiplying every fraction by the LCD 8 this will clear out the fractions

jimthompson5910 (jim_thompson5910):

7/8y-5/8=15/4 8*((7/8)y)-8(5/8)=8(15/4) 7y - 5 = 2*15 7y - 5 = 30 see what to do from here?

OpenStudy (zzelinski):

yeah, thanks so much!

OpenStudy (zzelinski):

I can figure out the others on my own, thanks

jimthompson5910 (jim_thompson5910):

ok glad to be of help